Listed below are the top 10 annual salaries​ (in millions of​ dollars) of TV personalities. Find the​ range, variance, and standard deviation for the sample data. Given that these are the top 10​ salaries, do we know anything about the variation of salaries of TV personalities in​ general?

42     41     39     33     21     18     15     11     10.6     9.3

The variance of the sample data is? ​(Round to two decimal places as​ needed.)
Answer #1
X (X - X̄) (X - X̄)²
42 18.01 324.36
41 17.01 289.34
39 15.01 225.30
33 9.01 81.18
21 -2.99 8.94
18 -5.99 35.88
15 -8.99 80.820
11 -12.99 168.740
10.6 -13.39 179.292
9.3 -14.69 215.796
X (X - X̄)²
total sum 239.9 1609.65
n 10 10

range=max-min =    42-9.3=   32.700
sample variance =    Σ(X - X̄)²/(n-1)=   1609.649/9=   178.850
sample std dev =   √ [ Σ(X - X̄)²/(n-1)] =   √(1609.649/9)=   13.373